OpenStudy (anonymous):

For those in Pre-Calc, can you help me with this problem? Solve each quadratic equation by the square root property: 3(x-4)^2=15

jimthompson5910 (jim_thompson5910):

\[\Large 3(x-4)^2=15\] \[\Large (x-4)^2=\frac{15}{3}\] \[\Large (x-4)^2=5\] \[\Large x-4=\pm\sqrt{5}\] \[\Large x=4\pm\sqrt{5}\]